# ModSec-AdvLearn: Countering Adversarial SQL Injections with Robust Machine Learning

### v3

@@ Abstract
Many Web Application Firewalls (WAFs) leverage the OWASP Core Rule Set (CRS) to block incoming malicious requests. The CRS consists of different sets of rules designed by domain experts to detect well-known web attack patterns. Both the set of rules to be used and the weights used to combine them are manually defined, yielding four different default configurations of the CRS. In this work, we focus on the detection of SQL injection (SQLi) attacks, and show that the manual configurations of the CRS typically yield a suboptimal trade-off between detection and false alarm rates. Furthermore, we show that these configurations are not robust to adversarial SQLi attacks, i.e., carefully-crafted attacks that iteratively refine the malicious SQLi payload by querying the target WAF to bypass detection. To overcome these limitations, we propose (i) using machine learning to automate the selection of the set of rules to be combined along with their weights, i.e., customizing the CRS configuration based on the monitored web services; and (ii) leveraging adversarial training to significantly improve its robustness to adversarial SQLi manipulations. Our experiments, conducted using the well-known open-source ModSecurity WAF equipped with the CRS rules, show that our approach, named ModSec-AdvLearn, can (i) increase the detection rate up to 30%, while retaining negligible false alarm rates and discarding up to 50% of the CRS rules; and (ii) improve robustness against adversarial SQLi attacks up to 85%, marking a significant stride toward designing more effective and robust WAFs. We release our open-source code at https://github.com/pralab/modsec-advlearn.
@@ I Introduction
While the domain knowledge poured in developing the CRS rules is extremely valuable, in this work we use the well-known ModSecurity WAF [9] equipped with the CRS rules to show that the heuristic choices made to select and combine such rules can lead to: (i) a suboptimal trade-off between detection rate and false alarms; and (ii) a substantial lack of robustness to adversarial SQLi attacks, i.e., functionality-preserving manipulations of the SQLi attack payload aimed to evade detection [10, 11]. To overcome these limitations, we then propose a novel robust machine learning approach to selecting and combining the CRS rules, named ModSec-AdvLearn, which is conceptually represented in Fig. 1 and detailed in Sect. III. This approach is built upon two main contributions. First, we propose using machine learning (ML) to automate both rule selection and weighing, adjusting the CRS configuration based on the traffic data collected from the monitored web services, and building on our preliminary findings in [11]. The underlying idea is to train a linear ML model using all the CRS rules as input features, aiming to improve the trade-off between detection and false alarm rates, by specializing the model to the specifics of the traffic data collected from the monitored applications. Furthermore, enforcing a sparse regularization during training enables the selection of an effective subset of rules as a by-product, avoiding the need for manual selection of the CRS rules in each PL. The second main contribution of this work is the definition of a novel adversarial training scheme to improve model robustness against adversarial SQLi attacks. To craft these attacks, we leverage WAF-A-MoLE [10], i.e., a black-box mutational fuzzer [12] that iteratively selects the best combination of random manipulations of SQLi payloads to reduce their probability of being detected by the targeted WAF. Nevertheless, for strengthening linear models, this may be not even required; one can equivalently use a specific regularizer to drastically reduce the computational overhead at training time, as shown in Sect. III-B.
@@
Through our extensive experimental analysis, conducted on two publicly-available datasets [10, 11], as reported in Sect. IV, we show that ModSec-AdvLearn is able to overcome the limitations of the current CRS, by detecting 30% attacks more while using much fewer rules. On top of this, ModSec-AdvLearn provides an unprecedented level of robustness against adversarial SQLi attacks, amounting to 85% more than the default CRS configurations. By deepening the investigation of this result, we discover that not only ModSec-AdvLearn relies more on rules that are not affected by adversarial SQLi attacks, but also that it gives more emphasis to additional rules that are triggered by some side-effect artifacts introduced by the adversarial SQLi manipulations.
@@
To conclude, let us remark that we are the first to demonstrate the effectiveness of adversarial training in the WAFs domain (specifically, for detecting SQLi attacks) when leveraging state-of-the-art input-space SQLi manipulations. This is completely different from other domains like image classifiers, where adversarial training is not sufficient to achieve a high level of robustness, and the effective mitigation of the risks presented by adversarial examples is still an open issue. We discuss these aspects along with related work in Sect. V, and the limitations of our approach as well as the corresponding future research directions in Sect. VI. We have also publicly released our code to foster reproducibility of our work. [^1] https://github.com/pralab/modsec-advlearn
@@ II-B The OWASP Core-Rule-Set (CRS) Project
Detection Rules. These are regular expressions (regex) that match specific byte patterns in requests. For instance, line 1 of 3 captures several patterns of comments commonly used in SQLi attacks such as ”;--” and ”-- ”. Each rule is denoted by a unique identifier (id in line 2), whose suffix also indicates the type of attack it is designed to identify (those starting with 942 target SQLi attacks [8]). Notable among configuration settings are the Paranoia Level (line 6) and Severity Level (line 7), which are explained below.
Paranoia Level. It defines the set of rules that are enabled to analyze the incoming HTTP requests [8]. The CRS includes four PLs (PL1 - PL4) and each rule is assigned to a specific PL; e.g., the rule in 3 belongs to PL1 (line 6). Moreover, rules are grouped together by PL in a nested way: setting a certain PL enables all the rules assigned to that PL, as well as those assigned to lower PLs. For instance, PL3 enables all the rules related to such PL, as well as those assigned to PL1 and PL2. Consequently, PL4 will enable all the rules.
Severity Level. Each CRS rule is heuristically given a severity level, i.e., a positive integer that quantifies how severe the corresponding attack could be [8]. The WAF applies the rules on each incoming request, and sums the severity levels of the firing rules. If the aggregated score exceeds a predefined threshold, the request is flagged as malicious. The CRS defines four severity levels: CRITICAL (5), ERROR (4), WARNING (3) and NOTICE (2); e.g., the severity level of the rule in 3 is CRITICAL (line 7), so it contributes to the aggregated score with a value of 5 (line 8).
@@ III-B ModSec-AdvLearn: Robustness against Adversarial SQLi
This finding sheds light on the role of the regularization term, showing that its choice should be based on the type of noise that affects the input data. In particular, it tells us that $\ell_{\infty}$ is the optimal regularizer if one wants to learn a robust model against $\ell_{1}$-norm (sparse) perturbations [22, 20].
@@
The given formulation corresponds to a linear programming problem, which can be solved using standard, off-the-shelf solvers, such as the simplex algorithm or interior-point techniques. We refer to this robust learning approach as Secure Support Vector Machine (SecSVM) in our experiments. [^5] Note that, even if we keep the same name of the approach proposed in [20], our SecSVM implementation is different, as we are neither using custom bounds on each feature value, nor any $\ell_{2}$ regularization.
Problem-space AT with WAF-A-MoLE. Let us now define a different approach to learn robust models directly against problem-space perturbations. The reason is that hardening models via feature-space AT or with ad-hoc regularization methods may provide an overly-pessimistic approach to feature manipulation that do not considers the specific constraint of realizable, semantic- and functionality-preserving SQLi attacks. In particular, practical SQLi manipulations may not enable switching on or off individual CRS rules independently, and may inadvertently trigger certain rules even if the attack aims to bypass the detection. Furthermore, such manipulations cannot be modeled as additive perturbations, and computing them typically requires inverting a complicated, non-differentiable feature extract step; in our case, this would amount to reversing the inner-workings of each CRS rule. For this reason, feasible SQLi attacks, as many other adversarial perturbations aimed to bypass ML models for cybersecurity-related tasks [21], cannot be typicaly optimized via gradient descent directly. To overcome these issues, we consider here a more general problem-space AT procedure, defined as:
@@
where $h(\boldsymbol{z},\boldsymbol{\delta})$ is a manipulation function that modifies the input SQL query $\boldsymbol{z}$ and returns a semantic-preserving SQL query $\boldsymbol{z}^{\prime}$, based on the choice of its input parameters $\boldsymbol{\delta}\in\Delta$. The set $\Delta$ constrains the input manipulations described in Table I to produce valid samples, e.g., picking only visible character when adding or re-writing comments, produce always-true or always-false conditions that do not change the original evaluation of the payload, or encoding integers in a specific base different from the original one.
@@ IV-A Experimental Setup
Datasets. We conduct our experiments using two datasets. The first one, WAF-A-MoLE Dataset [10], which consists of 393,629 malicious and 345,199 benign SQL queries. Benign samples were generated from a restricted SQL grammar, while attacks were generated using well-known web security testing tools such as SQLmap and OWASP ZAP [10]. The second one, ModSec-Learn Dataset [11], instead consists of 25,000 malicious SQLi payloads and 25,000 benign HTTP requests, based on real-world traffic. Legitimate samples were collected from the open-appsec dataset, [^6] https://github.com/openappsec/waf-comparison-project/tree/main/Data which contains samples from various real-world scenarios. Malicious samples were collected from multiple sources, and generated through security testing tools such as SQLmap, by executing it with different tampering scripts designed for payload obfuscation.
We divide each dataset into four subsets: training (train), test (test), adversarial training (train-adv), and adversarial test (test-adv). Table II shows the distribution of samples across the four subsets for each dataset: train contains an equal number of benign and SQLi queries, and it is used to train our target WAFs; test, disjoint from train, is used to evaluate the baseline performances of target WAFs, including vanilla ModSecurity, ModSec-Learn, and ModSec-AdvLearn, across different PLs; train-adv contains the same samples of train, but optimizes half of the SQLi queries using WAF-A-MoLE against the target WAF for problem-space AT; test-adv contains the same samples of test, but optimizes all the SQLi queries using WAF-A-MoLE to bypass the target WAF, i.e., to evaluate its robustness. We would like to clarify that, although using the same SQLi fuzzer (i.e., WAF-A-MoLE), the adversarial examples generated for building the adversarial training and test sets are different. They are independently optimized against each target model at test time, resulting in the application of different, optimal manipulation strategies. Hence, the two sets are independent, ensuring an unbiased evaluation.
@@
Machine Learning. We leverage scikit-learn v1.4.0 implementations of SVM (LinearSVC), LR, and RF to implement both ModSec-Learn and ModSec-AdvLearn. For the SVM and LR models, we experiment with both $\ell_{1}$ and $\ell_{2}$ regularizers. We implement SecSVM as described in Sect. III-B, using the linear programming solver provided by SciPy. We refer to it as ModSec-Learn SecSVM in the reported tables. The hyperparameters of each model are tuned via grid search, performing a 5-fold cross validation on the training set (train) to maximize the mean F1 score. In the case of ModSec-Learn, for the SVM and LR, we tune the regularization parameter $C\in\{10^{-3},10^{-2},10^{-1},0.5,1.0\}$. The best value found is typically $C=0.5$ for both models. For SecSVM, we tune the hyperparameter $t\in\{0.1,0.2,\ldots,1.0\}$. The best $t$ value is typically found to be $0.5$. The RF model is used with its default hyperparameters. In the case of ModSec-AdvLearn, adversarial training is applied only for PL4, given that the models trained on this PL demonstrated better performance on the test set (see Sect. IV-C). The hyperparameters are tuned using the same procedure detailed above, finding approximately the same best values, except for SecSVM, for which $t=1.0$ tends to provide better results.
Figure 2: ROC curves of vanilla ModSecurity (ModSec) and ModSec-Learn approaches (SVM, RF, LR, and SecSVM), evaluated on test (solid lines) and test-adv (dashed lines), for WAF-A-MoLE (top row) and ModSec-Learn (bottom row) datasets. Each curve reports the fraction of detected SQLi attacks (TPR) against the fraction of misclassified legitimate requests (FPR).
@@ IV-B Evaluation of ModSecurity
The first goal of our experimental analysis is understanding the detection capability of the vanilla ModSecurity. Rather than focusing only on CRS default values, we experiment with it over its entire configuration space, considering all the possible values for the PLs and the classification threshold. Hence, for each PL, we compute the Receiver-Operating-Characteristic (ROC) curve, which reports the detection rate, a.k.a. True Positive Rate (TPR, i.e., the fraction of correctly-detected malicious SQLi requests) against the False Positive Rate (FPR, i.e., the fraction of wrongly-classified legitimate requests) obtained by considering all possible classification threshold values. We report our findings with red lines in Fig. 2, while in Table III we extrapolate the TPR values at 1% FPR. We want to point out that, although the ROC curves in Fig. 2 already show the TPR for each possible operating point (i.e., the value of FPR), we report the results in Table III at 1% FPR because it is a reasonable value commonly adopted in the literature [23, 20]. We detail hereafter the key findings of our evaluations of ModSecurity against both the test set (test) and the adversarial test set (test-adv).
Evaluation on Clean Samples. We first test ModSecurity on the data we have gathered, and the results of this first evaluation are indicated with red solid lines in Fig. 2. The ROC curve of PL1 (default PL for ModSecurity) shows its inability to discriminate between benign and malicious SQL queries, with a TPR of 63.85% at a 1% False Positive Rate (FPR). The results for PL2 are the best among all the PLs, with a TPR of 66.82% at 1% FPR. The ROC curves for both PL3 and PL4 are almost identical, showing that PL4’s additional SQLi rules do not improve detection capabilities.
@@
TABLE III: TPR at 1% FPR for ModSecurity (ModSec) and ModSec-Learn (SVM, LR, SecSVM, and RF) evaluated on the baseline (test) and adversarial (test-adv) test sets. In addition, it also reports the evaluation of ModSec-AdvLearn (SVM, LR, SecSVM, and RF) for PL4. The column AR shows the number of active rules for each model. Results are shown for both datasets: WAF-A-MoLE and ModSec-Learn.

### v4

@@ Abstract
Many Web Application Firewalls (WAFs) leverage the OWASP Core Rule Set (CRS) to block incoming malicious requests. The CRS consists of different sets of rules designed by domain experts to detect well-known web attack patterns. Both the set of rules and the weights used to combine them are manually defined, yielding four different default configurations of the CRS. In this work, we focus on the detection of SQL injection (SQLi) attacks, and show that the manual configurations of the CRS typically yield a suboptimal trade-off between detection and false alarm rates. Furthermore, we show that these configurations are not robust to adversarial SQLi attacks, i.e., carefully-crafted attacks that iteratively refine the malicious SQLi payload by querying the target WAF to bypass detection. To overcome these limitations, we propose (i) using machine learning to automate the selection of the set of rules to be combined along with their weights, i.e., customizing the CRS configuration based on the monitored web services; and (ii) leveraging adversarial training to significantly improve its robustness to adversarial SQLi manipulations. Our experiments, conducted using the well-known open-source ModSecurity WAF equipped with the CRS rules, show that our approach, named ModSec-AdvLearn, can (i) increase the detection rate up to 30%, while retaining negligible false alarm rates and discarding up to 50% of the CRS rules; and (ii) improve robustness against adversarial SQLi attacks up to 85%, marking a significant stride toward designing more effective and robust WAFs. We release our open-source code at https://github.com/pralab/modsec-advlearn.
@@ I Introduction
While the domain knowledge poured in developing the CRS rules is extremely valuable, in this work we use the well-known ModSecurity WAF [9] equipped with the CRS rules to show that the heuristic choices made to select and combine such rules can lead to: (i) a suboptimal trade-off between detection rate and false alarms; and (ii) a substantial lack of robustness to adversarial SQLi attacks, i.e., functionality-preserving manipulations of the SQLi attack payload aimed to evade detection [10, 11]. To overcome these limitations, we then propose a novel robust machine learning approach to selecting and combining the CRS rules, named ModSec-AdvLearn, which is conceptually represented in Fig. 1 and detailed in Sect. III. This approach is built upon two main contributions. First, we propose using machine learning (ML) to automate both rule selection and weighting, adjusting the CRS configuration based on the traffic data collected from the monitored web services, and building on our preliminary findings in [11]. The underlying idea is to train a linear ML model using all the CRS rules as input features, aiming to improve the trade-off between detection and false alarm rates, by specializing the model to the specifics of the traffic data collected from the monitored applications. Furthermore, enforcing a sparse regularization during training enables the selection of an effective subset of rules as a by-product, avoiding the need for manual selection of the CRS rules in each PL. The second main contribution of this work is the definition of a novel adversarial training scheme to improve model robustness against adversarial SQLi attacks. To craft these attacks, we leverage WAF-A-MoLE [10], i.e., a black-box mutational fuzzer [12] that iteratively selects the best combination of random manipulations of SQLi payloads to reduce their probability of being detected by the targeted WAF. We also show in Sect. III-B that using $\ell_{\infty}$-norm regularization on a linear model yields equivalent robust solutions, avoiding the computational burden of optimizing attacks during training.
@@
Through our experiments, reported in Sect. IV and conducted on two publicly-available datasets [10, 11], we show that ModSec-AdvLearn overcomes the limitations of the current CRS, by detecting 30% attacks more with much fewer rules. ModSec-AdvLearn also provides an unprecedented level of robustness against adversarial SQLi attacks, i.e., 85% more than the default CRS configurations. By deepening the investigation of this result, we discover that ModSec-AdvLearn gives more emphasis to rules that are (i) less affected by adversarial SQLi attacks and (ii) accidentally triggered by side-effect artifacts introduced by the adversarial SQLi manipulations.
@@
To conclude, we remark that our work is the first to demonstrate the effectiveness of adversarial training in the WAFs domain (specifically, for detecting SQLi attacks) when leveraging state-of-the-art input-space SQLi manipulations. This is completely different from other domains like image classifiers, where adversarial training is not sufficient to achieve a high level of robustness, and the effective mitigation of the risks presented by adversarial examples is still an open issue. For this reason, we firmly believe that our work provides interesting and novel insights on how to design robust machine learning models for cybersecurity. We discuss these aspects along with related work in Sect. V, and the limitations of our approach as well as the corresponding future research directions in Sect. VI. We have also publicly released our code to foster reproducibility of our work. [^1] https://github.com/pralab/modsec-advlearn
@@ II-B The OWASP Core-Rule-Set (CRS) Project
Detection Rules. These are regular expressions (regex) that match specific byte patterns in requests. For instance, line 3 of 3 captures several patterns of comments commonly used in SQLi attacks such as ”;--” and ”-- ”. Each rule is denoted by a unique identifier (id in line 4), whose suffix also indicates the type of attack it is designed to identify (those starting with 942 target SQLi attacks [8]). Notable among configuration settings are the Paranoia Level (line 8) and Severity Level (line 9), which are explained below.
Paranoia Level. It defines the set of rules that are enabled to analyze the incoming HTTP requests [8]. The CRS includes four PLs (PL1 - PL4) and each rule is assigned to a specific PL; e.g., the rule in 3 belongs to PL1 (line 8). Moreover, rules are grouped together by PL in a nested way: setting a certain PL enables all the rules assigned to that PL, as well as those assigned to lower PLs. For instance, PL3 enables all the rules related to such PL, as well as those assigned to PL1 and PL2. Consequently, PL4 will enable all the rules.
Severity Level. Each CRS rule is heuristically given a severity level, i.e., a positive integer that quantifies how severe the corresponding attack could be [8]. The WAF applies the rules on each incoming request, and sums the severity levels of the firing rules. If the aggregated score exceeds a predefined threshold, the request is flagged as malicious. The CRS defines four severity levels: CRITICAL (5), ERROR (4), WARNING (3) and NOTICE (2); e.g., the severity level of the rule in 3 is CRITICAL (line 9), so it contributes to the aggregated score with a value of 5.
@@ III-B ModSec-AdvLearn: Robustness against Adversarial SQLi
This finding sheds light on the role of the regularization term, showing that its choice should be based on the type of noise that affects the input data. In particular, it tells us that $\ell_{\infty}$ is the optimal regularizer for training a robust model against $\ell_{1}$-norm (sparse) perturbations. It is also not difficult to see that, analogously, the standard $\ell_{2}$-norm SVM is optimal against $\ell_{2}$-norm (dense) perturbations [22, 20].
@@
The given formulation corresponds to a linear programming problem in its canonical form, which can be solved using standard, off-the-shelf solvers, such as the simplex algorithm or interior-point methods. In this work, the optimization problem is solved using the linear programming solver provided by the SciPy library [^5] https://docs.scipy.org/doc/scipy/reference/generated/scipy.optimize.linprog.html. We refer to this robust learning approach as Secure Support Vector Machine (SecSVM) in our experiments. [^6] Note that, even if we keep the same name of the approach proposed in [20], our SecSVM implementation is different, as we are neither using custom bounds on each feature value nor any $\ell_{2}$ regularization.
Problem-space AT with WAF-A-MoLE. Let us now define a different approach to learning robust models directly against problem-space perturbations. The reason is that hardening models via feature-space AT or with ad-hoc regularization methods may provide an overly pessimistic approach to feature manipulation that does not consider the specific constraint of realizable, semantic- and functionality-preserving SQLi attacks. In particular, practical SQLi manipulations may not enable switching on or off individual CRS rules independently, and may inadvertently trigger certain rules even if the attack aims to bypass the detection. Furthermore, such manipulations cannot be modeled as additive perturbations, and computing them typically requires inverting a complicated, non-differentiable feature extraction step; in our case, this would amount to reversing the inner workings of each CRS rule. For this reason, feasible SQLi attacks, as many other adversarial perturbations aimed to bypass ML models for cybersecurity-related tasks [21], cannot be typically optimized via gradient descent directly. To overcome these issues, we consider here a more general problem-space AT procedure, defined as:
@@
where $h(\boldsymbol{z},\boldsymbol{\delta})$ is a manipulation function that modifies the input SQL query $\boldsymbol{z}$ and returns a semantic-preserving SQL query $\boldsymbol{z}^{\prime}$, based on the choice of its input parameters $\boldsymbol{\delta}\in\Delta$. The set $\Delta$ constrains the input manipulations described in Table I to produce valid samples, e.g., picking only visible characters when adding or re-writing comments, produce always-true or always-false conditions that do not change the original evaluation of the payload, or encoding integers in a specific base different from the original one.
@@ IV-A Experimental Setup
Datasets. We conduct our experiments using two datasets. The first one, WAF-A-MoLE Dataset [10], which consists of 393,629 malicious and 345,199 benign SQL queries. Benign samples were generated from a restricted SQL grammar, while attacks were generated using well-known web security testing tools such as SQLmap and OWASP ZAP [10]. The second one, ModSec-Learn Dataset [11], instead consists of 25,000 malicious SQLi payloads and 25,000 benign HTTP requests, based on real-world traffic. Legitimate samples were collected from the open-appsec dataset, [^7] https://github.com/openappsec/waf-comparison-project/tree/main/Data, which contains samples from various real-world scenarios. Malicious samples were collected from multiple sources, and generated through security testing tools such as SQLmap, by executing it with different tampering scripts designed for payload obfuscation.
We divide each dataset into four subsets: training (train), test (test), adversarial training (train-adv), and adversarial test (test-adv). Table II shows the distribution of samples across the four subsets for each dataset: train contains an equal number of benign and SQLi queries, and it is used to train our target WAFs; test, disjoint from train, is used to evaluate the baseline performances of target WAFs, including vanilla ModSecurity, ModSec-Learn, and ModSec-AdvLearn, across different PLs; train-adv contains the same samples of train, but $50\%$ of the SQLi queries in the WAF-A-MoLE dataset and the $25\%$ of the SQLi queries in the ModSec-Learn dataset are optimized using WAF-A-MoLE against the target WAF for problem-space AT; test-adv contains the same samples of test, but optimizes all the SQLi queries using WAF-A-MoLE to bypass the target WAF, i.e., to evaluate its robustness. We would like to clarify that, although using the same SQLi fuzzer (i.e., WAF-A-MoLE), the adversarial examples generated for building the adversarial training and test sets are different. They are independently optimized against each target model at test time, resulting in the application of different, optimal manipulation strategies. Thus, the sets are independent, ensuring an unbiased evaluation.
@@
Machine Learning. We leverage scikit-learn v1.4.0 implementations of SVM (LinearSVC), LR, and RF to implement both ModSec-Learn and ModSec-AdvLearn. For the SVM and LR models, we experiment with both $\ell_{1}$ and $\ell_{2}$ regularizers. We implement SecSVM as described in Sect. III-B, using the linear programming solver provided by SciPy. We refer to it as ModSec-Learn SecSVM in the reported tables. The hyperparameters of each model are tuned via grid search, performing a 5-fold cross validation on the training set (train) to maximize the mean F1 score. In the case of ModSec-Learn, for the SVM and LR, we tune the regularization parameter $C\in\{10^{-3},10^{-2},10^{-1},0.5,1.0\}$. The best value found is typically $C=0.5$ for both models. For SecSVM, we tune the hyperparameter $t\in\{0.1,0.2,\ldots,1.0\}$. The best $t$ value is typically found to be $0.5$. The RF model is used with its default hyperparameters. In the case of ModSec-AdvLearn, adversarial training is applied only for PL4, given that the models trained on this PL demonstrated better performance on the test set (see Sect. IV-C). The hyperparameters are tuned using the same procedure described in this paragraph, finding approximately the same best values, except for SecSVM, for which $t=1.0$ yields better results.
Figure 2: ROC curves of vanilla ModSecurity (ModSec) and ModSec-Learn approaches (SVM, RF, LR, and SecSVM), evaluated on test (solid lines) and test-adv (dashed lines), for WAF-A-MoLE (first two rows) and ModSec-Learn (last two rows) datasets. Each curve reports the fraction of detected SQLi attacks against the fraction of misclassified legitimate requests.
@@ IV-B Evaluation of ModSecurity
The first goal of our experimental analysis is to understand the detection capability of the vanilla ModSecurity. Rather than focusing only on CRS default values, we experiment with it over its entire configuration space, considering all the possible values for the PLs and the classification threshold. Hence, for each PL, we compute the Receiver-Operating-Characteristic (ROC) curve, which reports the detection rate, a.k.a. True Positive Rate (TPR, i.e., the fraction of correctly-detected malicious SQLi requests) against the False Positive Rate (FPR, i.e., the fraction of wrongly-classified legitimate requests) obtained by considering all possible classification threshold values. We report our findings with red lines in Fig. 2, while in Table III we extrapolate the TPR values at 1% FPR. We want to point out that, although the ROC curves in Fig. 2 already show the TPR for each possible operating point (i.e., the value of FPR), we report the results in Table III at 1% FPR because it is a reasonable value commonly adopted in the literature [23, 20]. We detail hereafter the key findings of our evaluations of ModSecurity against both the test set (test) and the adversarial test set (test-adv).
Evaluation on Clean Samples. We first test ModSecurity on the data we have gathered, and the results of this first evaluation are indicated with red solid lines in Fig. 2. The ROC curve of PL1 (default PL for ModSecurity) shows its inability to discriminate between benign and malicious SQL queries, with a TPR of 63.85% at a 1% False Positive Rate (FPR). The results for PL2 are the best among all PLs, with a TPR of 66.82% at 1% FPR. The ROC curves for both PL3 and PL4 are almost identical, as PL4 has only two active rules more than PL3, which do not even improve its detection rate.
@@
TABLE III: TPR at 1% FPR evaluated on the baseline/adversarial (test/test-adv) test sets for ModSecurity (ModSec); ModSec-Learn (SVM, LR, SecSVM, and RF); and ModSec-AdvLearn (SVM, LR, SecSVM, and RF) using PL4. We also report the number of active rules (AR) for each model. Results are shown for both datasets: WAF-A-MoLE and ModSec-Learn.